# Improving Reward Models with Synthetic Critiques

### v1

@@ Abstract
Reward models (RM) play a critical role in aligning language models through the process of reinforcement learning from human feedback. RMs are trained to predict a score reflecting human preference, which requires significant time and cost for human annotation. Additionally, RMs tend to quickly overfit on superficial features in the training set, hindering their generalization performance on unseen distributions. We propose a novel approach using synthetic natural language critiques generated by large language models to provide additional feedback, evaluating aspects such as instruction following, correctness, and style. This offers richer signals and more robust features for RMs to assess and score on. We demonstrate that high-quality critiques improve the performance and data efficiency of RMs initialized from different pretrained models. Conversely, we also show that low-quality critiques negatively impact performance. Furthermore, incorporating critiques enhances the interpretability and robustness of RM training.
@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has become a popular technique for aligning large language models (LLMs) with human preferences, underpinning the success of recent state-of-the-art LLMs (Ouyang et al., 2022; Bai et al., 2022; Achiam et al., 2023). RLHF involves two main steps. First, a reward model (RM) is trained to generate scalar rewards based on human preferences. In the second step, reinforcement learning algorithms, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), are used to optimize language models by maximizing the rewards predicted by the trained RMs. The reward model plays a crucial role in the RLHF process by serving as a proxy for human preferences, predicting whether a user is likely to prefer one piece of text over another. The RM is built on preference data collected from human annotators. Given an input instruction, it compares two completions and rates the better one with a higher score. Underscoring the importance of the reward model, Llama 2 (Touvron et al., 2023a) used 1 million such binary preference datapoints to train their reward model.
@@
Simultaneously, the prevailing approach to training preference-based reward models presents several challenges: understanding how RLHF processes align with human preferences is challenging due to a lack of explainability and interpretability in RM training. Preference scores annotated by humans can be noisy, subjective and open to biases (Hosking et al., 2023; Kirk et al., 2024). Additionally, it is difficult to determine whether the RMs are fitting to superficial features or learning the actual preferences expressed by humans in the data, and consequently how such values are internalized by the LLMs during the second preference tuning stage (Casper et al., 2023; Marks et al., 2024; Qiu et al., 2024). This process may also introduce biases, for example, towards longer responses (Singhal et al., 2023).
Reward models can also be very brittle – coupled with them picking up on data artefacts, this can lead to over-optimization towards the training distribution which leads to inconsistencies between RM predictions and human preferences at inference time. Challenges also include generalization to distribution shifts, as well as susceptibility to adversarial attacks (Shen et al., 2023a; Gao et al., 2023; Coste et al., 2023). A final drawback is that training reward models requires training data, of a different kind than the one used to train generative text models. Collecting these human preference annotations can be costly and laborious (Cui et al., 2023). As a consequence, existing preferences datasets are small in scale (Wu et al., 2023), and limited on specific tasks (Stiennon et al., 2022).
In this paper, we propose using model-generated natural language critiques to overcome these limitations. We illustrate our proposal in Figure 1. In step (1), we prompt LLMs to generate synthetic critiques for each point-wise prompt-completion pair in the preference data. These critiques evaluate responses across multiple dimensions, including correctness and adherence to instructions. In step (2), we train RMs conditioned on these critiques to predict scalar rewards and evaluate their performance on a test set augmented with critiques. As synthetic critiques can be efficiently generated using open-source models, our method is both accessible and cost-effective.
@@ 2. RQ2:
    What impact do critiques have on RM performance and data efficiency as training scales up?
Our experiments show that using such critiques improves RM performance, especially in the low-resource data regime. In particular, we show that a high-quality model-generated critique is roughly equivalent to 40 vanilla preference pairs, thus paving the way for more efficient use of annotation budgets. The improvement is particularly striking in subsets that are adversarially created, showing that critiques can limit the impact of superficial features on determining the final reward scores.
@@ 2 Related Work
RLHF usually involves training a reward model on preference data collected from annotators that is used for learning a policy using RL algorithms like Reinforce (Ahmadian et al., 2024) or PPO (Schulman et al., 2017). Different methods have been proposed for solving the reward modeling tasks. One of the most common ways to approach reward modeling is to apply the Bradley-Terry model (Bradley and Terry, 1952) for predicting human preference. Rafailov et al. (2024) proposed Direct Preference Optimization, where instead of learning a separate reward model, the LLM can be used as a proxy reward model. Note that even these offline approaches require preference data, though without the need to train a separate reward model. Lastly, there are many efforts that use generative LLMs as reward models by prompting the systems to choose a preference (Li et al., 2023b, 2024; Zheng et al., 2023; Verga et al., 2024) or directly output a score (Cui et al., 2023).
@@
Due to the central role that reward models play in the RLHF process, there is an increasing interest for systematically understanding their behaviour. Lambert et al. (2024) introduced RewardBench, a benchmark dataset for evaluating reward models that contains pairs of completions with subtle but meaningful reasons of why one completion is preferred over the other. RewardBench showed that while reward models are able to perform well on general chat, they struggle with more challenging reasoning examples. Apart from general performance, reward modeling is affected by issues such as length bias (Singhal et al., 2023; Shen et al., 2023b) and lack of interpretability (Clymer et al., 2023). Moreover, these undesirable biases also mislead human annotators (Hosking et al., 2023).
@@
Critiques have emerged as a method for improving interpretability during the model development process and robustness against adversarial examples (Wang et al., 2023a). These rationales have been proven to play an essential role during LLM training (Scheurer et al., 2023; Wu et al., 2023) and evaluation (Zheng et al., 2023; Li et al., 2023a). Recent works explore using critiques for self-improvement by refining their outputs into better ones (Madaan et al., 2024; Gou et al., 2023; Ye et al., 2023). Additionally, Saunders et al. (2022) demonstrate how critiques are helpful for human annotators in identifying flaws that they might have otherwise missed, while Yuan et al. (2024) analyze the use of critiques to provide feedback to human writers.
@@
More similar to our work, Zeng et al. (2023) find that when using language models as evaluators, prompting them with a rationale and rubric improves their evaluator accuracy. Auto-J (Li et al., 2023a) extends this idea by training a generative LLM judge that is able to rate responses and generate critiques, improving the overall accuracy of the model. Despite critiques showing positive results in a wide variety of applications, their success heavily relies on their quality (Sun et al., 2024). In order to measure critiques quality, Lan et al. (2024) propose CriticBench, a benchmark designed to assess four dimensions of critique ability in LLMs: feedback, comparison, refinement, and meta-feedback.
@@ 3.1 Synthetic Critique Generation with LLMs
In the first step, we generate synthetic critiques for the preference data $\mathcal{D}$ used to train the RMs. We prompt LLMs to generate natural language critiques as follows: Given a preference data pair $(x,y_{+},y_{-})$, where $x$ is a prompt and $(y_{+},y_{-})$ are the chosen and rejected completions, we prompt the LLM to generate point-wise critiques $c$ for each completion $y$. Specifically, for each chosen pair ($x$, $y$), where $y$ can be either a chosen or rejected completion, we generate a critique $c$. This way, we obtain a new critique-augmented training dataset with triplets $(x,y,c)$. We design prompts for LLMs to generate synthetic critiques that evaluate the prompt-completion pair based on how effectively the completion fulfills the prompt requirements in aspects such as instruction-following, truthfulness, and helpfulness. An example of critique is shown in Figure 1. For details of the prompt templates see Appendix A.
@@ 4.1 Datasets
For RM training, we collect a human-preference dataset that consists of 5k examples of open-ended, multi-turn conversations between a user and a chatbot. Each dataset entry includes a prompt or input instruction and two corresponding completions, with a human-annotated label indicating the better option. A preference rating is also given on a scale of three points (slightly better, better, significantly better). For more details of training data see Appendix B. During inference, for each prompt paired with two completions from the test set, we use the trained RMs to assign a reward score for both the chosen and rejected completion. We use test accuracy as a metric to assess the performance of the RM, calculated based on the proportion of instances where the chosen completion receives a higher score than the rejected one. [^2] We use accuracy on test prompts as a metric for evaluating RMs, under the premise that higher RM test accuracy correlates with enhanced downstream performance of LLMs trained with these RMs in RLHF.
We evaluate the RMs on a variety of benchmark datasets that target a broad set of capabilities, including chat, instruction following, coding, and safety. A summary of evaluation dataset description is shown in Table 1. Specifically, we evaluate on Rewardbench (Lambert et al., 2024), a dataset of prompt completion pairs that benchmarks how RMs perform on challenging, structured and out-of-distribution queries. Rewardbench contains four categories with a total of 2,985 prompts from different subsets: Chat (358 prompt in total) contains prompts from AlpacaEval (Li et al., 2023b) and MTBench (Zheng et al., 2023) that test RM’s basic capability to distinguish correct responses in open-ended chat. Chat Hard (456) contains questions from LLMBar (Zeng et al., 2023) and MTBench that stress test RM’s ability to understand trick questions and subtly different instruction responses, where the dispreferred output has appealing superficial qualities that challenge LLM-as-a-judge evaluators. Safety (740) contains pairs from XSTest (Röttger et al., 2023) and Do-Not-Answer (Wang et al., 2023b) that test the model’s ability to refuse dangerous content and avoid incorrect refusals. Reasoning (1431) evaluates code and reasoning abilities. Code prompts from HumanEvalPack (Muennighoff et al., 2023) have correct and buggy code as chosen as rejected completions, and reasoning prompts are from PRM800k (Lightman et al., 2023). We also compute a Avg. Score by averaging the four categories with the same weighting in RewardBench.
@@ 4.2 LLM Critique Generator
To examine the effects of critiques, we select a range of LLMs with varying architectures, model sizes and training data for critique generation. We set up a pool of 6 models: (1) LLaMA2-7B-Chat (Touvron et al., 2023a), (2) LLaMA2-70B-Chat (Touvron et al., 2023a), (3) Mixtral-8x7B-Instruct (Jiang et al., 2024), (4) Command R, [^3] https://huggingface.co/CohereForAI/c4ai-command-r-v01 a 35B-parameter model, (5) Command R+ [^4] https://huggingface.co/CohereForAI/c4ai-command-r-plus (103B), and (6) GPT4-Turbo [^5] To be precise, we used gpt-4-0125-preview. (Achiam et al., 2023). Given a set of training and test preference data, we prompt each of the models to generate a set of critiques for both sets. For each experiment, we ensure that the train and test data match in that they are enriched by critiques generated by the same model. For baseline, we compare with No-Critiques, where we train and evaluate the RM on sets without critiques.
@@ 5 Evaluation Results
Cohere May 2024 0.964 0.713 0.927 0.977 -
C4AI Command R+ (Generative) 0.951 0.576 0.556 0.704 -
Table 2: Test accuracy on RewardBench and PandaLM for RMs starting from three pretrained model checkpoints and enriched with various model critiques. Best results are bolded, and second-best are underlined. State-of-the-art results are taken from the RewardBench leaderboard at the time of writing (May 2024).
@@ 5.1 Do synthetic critiques improve RM performance?
Table 2 summarizes the evaluation results on the four subsets in RewardBench, and PandaLM. Figure 2 displays RewardBench Avg. Score. In most cases, adding critiques improves RM test accuracy compared to No-Critiques baseline, illustrating the effectiveness of synthetic critiques.
@@
We observe that critique quality has an impact on RM test performances. Specifically, across the base pretrained models, stronger models such as GPT4-Turbo consistently provide the highest Avg. Score for both datasets, with a score increase of 0.16 and 0.14 on RewardBench for RMs starting from LLaMA-7B Base and Command-35B Base respectively. This is followed by Command R+, with an increase of 0.1 and 0.062 respectively. On the other hand, weaker critiques generated from a smaller model such as LLaMa2-7B-Chat have a detrimental effect on test accuracy on LLaMA-7B Base. Similar to Saunders et al. (2022) and Sun et al. (2024), we observe a positive correlation between the effect of critiques on test performance and critique model sizes, as shown in Figure 7. For more quantitative and qualitative analysis of critique quality with metacritique scores (Sun et al., 2024), see Appendix E.
@@ 5.2 What impacts do critiques have on RMs as training scales up?
We initialize RMs from the 35B Command R base model and generate synthetic critiques using Command R+. We use RewardBench Avg. Score for evaluation, as illustrated in Figure 3. We observe that on RewardBench, increasing the number of training examples generally increases test accuracy, both with and without critiques. Using synthetic critiques enhances data efficiency, as demonstrated by the greater accuracy gains of Command R+ Critiques over No-Critiques, particularly in low-data settings (5k-50k). The marginal benefit of critiques decreases in the presence of abundant data, as both RMs reach similar test performance at  200k training examples.
@@
How many No-Critiques examples Is a Critique Worth?
To illustrate the effectiveness of strong critiques, we plot the RewardBench score from Table 2 for RMs trained with 5k synthetic critiques using Command R, LLaMA2-70B-Chat, Command R+, and GPT4-Turbo as dashed lines in the background of Figure 3. We observe that 5k Command R+ Critiques achieve a test accuracy comparable to approximately 90k No-Critiques. Similarly, 5k Command R and 5k LLaMA2-70B-Chat equate to roughly 62k No-Critiques, while 5k GPT4-Turbo surpass even 200k No-Critiques. For a more detailed breakdown of test accuracy over different tasks under RewardBench with scaling see Appendix D. Interestingly, we observe that critiques consistently enhance performance on more difficult tasks, such as Reasoning, across increasing training data sizes. We hypothesize that this is due to the Chain-of-Thought style feedback within critiques, which provides additional context for RMs to condition on when predicting a scalar score. We leave it as future work to explore the impact of critiques on LLM reasoning abilities when preference-tuned with such critique-enhanced RMs.
5.3 Analysis
Finegrained Analysis
We find that the accuracy on the Chat subset of RewardBench is consistently low in Table  2 (often below random choice). To analyze this we repeat the scaling experiment, but evaluate only on this subset. The results in Figure 4(a) show how there is a significant improvement after the RM has seen 10k pairs, obtaining very high scores (0.91) and arguably saturating this task. Compare this to Figure 4(b) on Reasoning – a harder task – where using critiques outperforms the baseline (by more than 10 points) and where they only converge with 20k training examples. In addition, training with critiques leads improved robustness against adversarial examples on Chat Hard, and the effect is persistant across training data sizes. For more analysis see Appendix D.
@@ 5.3 Analysis
Comparison with Generative Judge RM
To validate the capacity of the trained RMs, we compare our trained classifier-based RMs with inference-time generative RMs (Li et al., 2023a) conditioned on critiques. We compare the classifier-based RMs starting from Command R-35B with a generative RM, as shown in the bottom section of Table 2 and the in background of Fig. 2. Specifically, we prompt Command R+ to directly output a score between 0 and 10 for each point-wise prompt-completion pair, and compute the proportion that is correct if the generated score is higher on the preferred completion. We call this setting Gen-Command R+. In Gen-Command R+ (+Critiques), we prompt for a critique before outputting a score. We use the same prompt template for classifier-based RMs (shown in Appendix A). We observe that similar to classifier-based models, conditioning on critiques leads to a higher test score than No-Critiques, agreeing with the observations in Wang et al. (2023a). We observe that the classifier-based RMs overwhelmingly outperform the generative judge baseline [^6] We include the performance of another generative RM from the public RewardBench Leaderboard, C4AI Command R+, in Table 2. Our generative RM accuracy differs from this model as we use point-wise evaluation, whereas C4AI Command R+ use pair-wise evaluation leading to higher scores. In pair-wise evaluation both completions are seen, arguably making the task easier., validating their effectiveness. While the generative judge performs strongly on Chat, comparably on Safety, it performs poorly in Reasoning, followed by Chat Hard. Upon closer inspection on Reasoning, we see that many of the mistakes for Gen-Command R+ (score of 0.060.06~{}0.060.06) are due to the generative judge assigning the same score to both preferred and worse, leading to a tie. We hypothesize that this is because without preference training, completion pairs from the more difficult Reasoning subset appear similar to the pretrained checkpoint, which is not sensitive to their differences. However, when conditioning on critiques, Gen-Command R+ (+Critiques) spreads out the reward scores on those ties, leading to a $0.3$ increase (score of $0.3685$).

### v2

@@ Abstract
Reward models (RMs) play a critical role in aligning language models through the process of reinforcement learning from human feedback. RMs are trained to predict a score reflecting human preference, which requires significant time and cost for human annotation. Additionally, RMs tend to quickly overfit on superficial features in the training set, hindering their generalization performance on unseen distributions. We propose a novel approach using synthetic natural language critiques generated by large language models to provide additional feedback, evaluating aspects such as instruction following, correctness, and style. This offers richer signals and more robust features for RMs to assess and score on. We demonstrate that high-quality critiques improve the performance and data efficiency of RMs initialized from different pretrained models, reducing the reliance on costly human annotations. Furthermore, incorporating critiques improves both the interpretability and robustness of RM training.
@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has emerged as a popular technique for aligning large language models (LLMs) with human preferences, underpinning the success of recent state-of-the-art LLMs (Ouyang et al., 2022; Bai et al., 2022; Achiam et al., 2023). RLHF typically involves two stages. First, a reward model (RM) is trained to generate scalar rewards based on human preferences. In the second stage, reinforcement learning algorithms are employed to optimize language models by maximizing the rewards predicted by the trained RMs. The reward model plays a crucial role in the RLHF process, acting as a proxy for human assessment. In preference-based algorithms, it is used to estimate whether a user is likely to favor one text over another, built upon preference data collected from human annotators. Given an input instruction, the RM compares two completions and assigns a higher score to the preferred option. Highlighting the significance of the reward model, Llama 2 (Touvron et al., 2023a) utilized an extensive dataset of 1 million such binary preference datapoints to train their reward model, emphasizing its crucial role in achieving alignment with human preferences.
@@
Simultaneously, the prevailing approach to training preference-based reward models presents several challenges: understanding how RLHF processes align with human preferences is challenging due to a lack of explainability and interpretability in RM training. Preference scores annotated by humans can be noisy, subjective and prone to biases (Hosking et al., 2023; Kirk et al., 2024). Additionally, it is difficult to determine whether the RMs are fitting to superficial features or learning the actual preferences expressed by humans in the data, and consequently how such values are internalized by the LLMs during the second preference tuning stage (Casper et al., 2023; Marks et al., 2024; Qiu et al., 2024). This process may also introduce biases, for example, favoring longer responses (Singhal et al., 2023).
Reward models can also be brittle – coupled with their tendency to pick up on data artefacts, this can lead to over-optimization towards the training distribution, resulting in inconsistencies between RM predictions and human preferences at inference time. Challenges also include generalization to distribution shifts, as well as susceptibility to adversarial attacks (Shen et al., 2023a; Gao et al., 2023; Coste et al., 2023). A final drawback is that training reward models requires training data, of a different kind than the one used to train generative text models. Collecting these human preference annotations can be costly and laborious (Cui et al., 2023). Consequently, existing preferences datasets are often small in scale (Wu et al., 2023), and limited on specific tasks (Stiennon et al., 2022).
In this paper, we propose to use model-generated natural language critiques to overcome these limitations, where the critiques explicitly reason about the quality of a response. We illustrate our proposal in Fig. 1. In step (1), we prompt LLMs to generate synthetic critiques for each point-wise prompt-completion pair in the preference data. These critiques evaluate responses across multiple dimensions, including correctness and adherence to instructions, in a manner similar to Chain-of-Thought (CoT) reasoning (Wei et al., 2022). In step (2), we train RMs conditioned on these critiques to predict scalar rewards and evaluate their performance on a test set augmented with critiques. We hypothesize that critiques, by explicitly reasoning about the quality of responses and extracting textual features, can address the challenges posed by implicit and potentially ambiguous human annotations. We also aim to explore whether these critiques, by highlighting meaningful and task-relevant aspects, can enhance RM training efficiency and robustness.
@@ 2. RQ2:
    How do critiques impact RM performance and data efficiency as training scales up?
Our experiments show that using synthetic critiques improves RM performance, particularly in the low-resource data regime. Notably, we show that a high-quality model-generated critique is roughly equivalent to 40 vanilla preference pairs, thus paving the way for more efficient use of annotation budgets. Additionally, we observe that critiques yield greater performance improvements on tasks requiring reasoning and adversarial skills, across various training data sizes. As synthetic critiques can be efficiently generated using open-source models, our method is both accessible and cost-effective.
@@ 2 Related Work
RLHF usually involves training a reward model on preference data collected from annotators that is used for learning a policy using RL algorithms like RLOO (Ahmadian et al., 2024) or PPO (Schulman et al., 2017). Different methods have been proposed for solving the reward modeling tasks. One of the most common ways to approach reward modeling is to apply the Bradley-Terry model (Bradley and Terry, 1952) for predicting human preference. Rafailov et al. (2024) proposed Direct Preference Optimization, where instead of learning a separate reward model, the LLM can be used as a proxy reward model. Note that even these offline approaches require preference data, though without the need to train a separate reward model. Lastly, there are many efforts that use generative LLMs as reward models by prompting the systems to choose a preference (Li et al., 2023b, 2024; Zheng et al., 2023; Verga et al., 2024) or directly output a score (Cui et al., 2023).
@@
Due to the central role that reward models play in the RLHF process, there is an increasing interest for systematically understanding their behaviour. Lambert et al. (2024) introduced RewardBench, a benchmark dataset for evaluating reward models that contains pairs of completions with subtle but meaningful reasons of why one completion is preferred over the other. RewardBench demonstrated that while reward models are able to perform well on general chat, they struggle with more challenging reasoning examples. Apart from general performance, reward modeling is affected by issues such as length bias (Singhal et al., 2023; Shen et al., 2023b) and lack of interpretability (Clymer et al., 2023). Moreover, these undesirable biases also mislead human annotators (Hosking et al., 2023).
@@
Critiques have emerged as a method for improving interpretability during the model development process and robustness against adversarial examples (Wang et al., 2023a). These rationales have been proven to play an essential role during LLM training (Scheurer et al., 2023; Wu et al., 2023) and evaluation (Zheng et al., 2023; Li et al., 2023a). Recent works explore using critiques for self-improvement by refining their outputs into better ones (Madaan et al., 2024; Gou et al., 2023; Ye et al., 2023). Additionally, Saunders et al. (2022) demonstrate how critiques are helpful for human annotators in identifying flaws that they might have otherwise missed, while Yuan et al. (2024) analyze the use of critiques to provide feedback to humans.
@@
More similar to our work, Zeng et al. (2023) find that when using language models as evaluators, prompting them with a rationale and rubric improves their evaluator accuracy. Auto-J (Li et al., 2023a) extends this idea by training a generative LLM judge that is able to rate responses and generate critiques, improving the overall accuracy of the model. Concurrent works that explore using critiques for training RMs include Ankner et al. (2024). Despite critiques showing positive results in a wide variety of applications, their success heavily relies on their quality (Sun et al., 2024). In order to measure critiques quality, Lan et al. (2024) propose CriticBench, a benchmark designed to assess four dimensions of critique ability in LLMs: feedback, comparison, refinement, and meta-feedback.
@@ 3.1 Synthetic Critique Generation with LLMs
In the first step, we generate synthetic critiques for the preference data $\mathcal{D}$ used to train the RMs. We prompt LLMs to generate critiques as follows: Given a preference data pair $(x,y_{+},y_{-})$, where $x$ is a prompt and $(y_{+},y_{-})$ are the chosen and rejected completions, we prompt the LLM to generate point-wise critiques $c$ for each completion $y$. Specifically, for each chosen pair ($x$, $y$), where $y$ can be either a chosen or rejected completion, we generate a critique $c$. This way, we obtain a new critique-augmented training dataset with triplets $(x,y,c)$. We used the critique prompt from Cui et al. (2023), designed to optimize feedback quality, to generate synthetic critiques $c$ that evaluate how well the completion meets prompt requirements, including instruction-following, truthfulness, and helpfulness. For details of the prompt template see App. A. An example of critique is shown in Fig. 1.
@@ 4.1 Datasets
For RM training, we collect a human-preference dataset that consists of 5k examples of open-ended, multi-turn conversations between a user and a chatbot. Each dataset entry includes a prompt or input instruction and two corresponding completions, with a human-annotated label indicating the better option. A preference rating is also given on a scale of three points (slightly better, better, significantly better). For more details of training data see App. B. During inference, for each prompt paired with two completions from the test set, we use the trained RMs to assign a reward score for both the chosen and rejected completion. We use test accuracy as a metric to assess the performance of the RM, calculated based on the proportion of instances where the chosen completion receives a higher score than the rejected one. [^2] We use accuracy on test prompts as a metric for evaluating RMs, under the premise that higher RM test accuracy correlates with enhanced downstream performance of LLMs trained with these RMs in RLHF.
We evaluate the RMs on a variety of benchmark datasets that target a broad set of capabilities, including chat, instruction following, coding, and safety. A summary of the evaluation dataset is shown in Table 1. Specifically, we evaluate on Rewardbench (Lambert et al., 2024), a dataset of prompt completion pairs that benchmarks how RMs perform on challenging, structured and out-of-distribution queries. Rewardbench contains four categories with a total of 2,985 prompts from different subsets: Chat (358 prompt in total) contains prompts from AlpacaEval (Li et al., 2023b) and MTBench (Zheng et al., 2023) that test RM’s basic capability to distinguish correct responses in open-ended chat. Chat Hard (456) contains adversarial prompts from LLMBar (Zeng et al., 2023) and MTBench that stress test RM’s ability to understand trick questions and subtly different instruction responses, where the dispreferred output has appealing superficial qualities that challenge LLM-as-a-judge evaluators. Safety (740) contains pairs from XSTest (Röttger et al., 2023) and Do-Not-Answer (Wang et al., 2023b) that test the model’s ability to refuse dangerous content and avoid incorrect refusals. Reasoning (1431) evaluates code and reasoning abilities. Code prompts from HumanEvalPack (Muennighoff et al., 2023) have correct and buggy code as chosen as rejected completions, and reasoning prompts are from PRM800k (Lightman et al., 2023). We also compute a Avg. Score by averaging the four categories with the same weighting as in RewardBench.
@@ 4.2 LLM Critique Generator
To examine the effects of critiques, we select a range of LLMs with varying architectures, model sizes and training data for critique generation. We set up a pool of 6 models: (1) LLaMA2-7B-Chat (Touvron et al., 2023a), (2) LLaMA2-70B-Chat (Touvron et al., 2023a), (3) Mixtral-8x7B-Instruct (Jiang et al., 2024), (4) Command R,333huggingface.co/CohereForAI/c4ai-command-r-v01 a 35B-parameter model, (5) Command R+444huggingface.co/CohereForAI/c4ai-command-r-plus (103B), and (6) GPT4-Turbo [^5] To be precise, we used gpt-4-0125-preview. (Achiam et al., 2023). Given a set of training and test preference data, we prompt each of the models to generate a set of critiques for both sets. For each experiment, we ensure that the train and test data match in that they are enriched by critiques generated by the same model. For baseline, we compare with No-Critiques, where we train and evaluate the RM on sets without critiques.
@@ 5 Evaluation Results
TextEval-Llama3.1-70B 0.941 0.901 0.932 0.964
Llama-3.1-Nemotron-70B-Reward 0.975 0.857 0.951 0.981
Table 2: Test accuracy on RewardBench and PandaLM for RMs starting from three pretrained model checkpoints and enriched with various model critiques. Best results are bolded, and second-best are underlined. State-of-the-art results are taken from the RewardBench leaderboard (Oct 2024).
@@ 5.1 Do synthetic critiques improve RM performance?
Table 2 summarizes the evaluation results on the four subsets in RewardBench, and PandaLM. Fig. 2 displays RewardBench Avg. Score. In most cases, adding critiques improves RM test accuracy compared to No-Critiques baseline, illustrating the effectiveness of synthetic critiques.
@@
We observe that critique quality has an impact on RM test performances. Specifically, across the base pretrained models, stronger models such as GPT4-Turbo consistently provide the highest Avg. Score for both datasets, with a score increase of 0.16 and 0.14 on RewardBench for RMs starting from LLaMA-7B Base and Command-35B Base respectively. This is followed by Command R+, with an increase of 0.1 and 0.062 respectively. On the other hand, weaker critiques generated from a smaller model such as LLaMa2-7B-Chat have a detrimental effect on test accuracy on LLaMA-7B Base. Similar to Saunders et al. (2022) and Sun et al. (2024), we observe a positive correlation between the effect of critiques on test performance and critique model sizes, as shown in Fig. 7. For more quantitative and qualitative analysis of critique quality with metacritique scores (Sun et al., 2024), see App. E.
@@ 5.2 What impacts do critiques have on RMs as training scales up?
We initialize RMs from the 35B Command R base model and generate synthetic critiques using Command R+. We use RewardBench Avg. Score for evaluation, as illustrated in Fig. 3. We observe that on RewardBench, increasing the number of training examples generally increases test accuracy, both with and without critiques. Using synthetic critiques enhances data efficiency, as demonstrated by the greater accuracy gains of Command R+ Critiques over No-Critiques, particularly in low-data settings (5k-50k). The benefit of critiques becomes less pronounced with abundant data, as both RMs achieve comparable test performance around 200k training examples. However, when critique quality is high, we continue to observe a substantial performance boost even in abundant-data settings: the performance gap remains evident at 200k examples, with just 5k strong critique models (dashed line of GPT-4 in Fig. 3).
@@
How many No-Critiques examples Is a Critique Worth?
To illustrate the effectiveness of strong critiques, we plot the RewardBench score from Table 2 for RMs trained with 5k synthetic critiques using Command R, LLaMA2-70B-Chat, Command R+, and GPT4-Turbo as dashed lines in the background of Fig. 3. We observe that 5k Command R+ Critiques achieve a test accuracy comparable to approximately 90k No-Critiques. Similarly, 5k Command R and 5k LLaMA2-70B-Chat equate to roughly 62k No-Critiques, while 5k GPT4-Turbo surpass even 200k No-Critiques.
5.3 Analysis
Finegrained Analysis
We observe that the benefit of critiques varies by task type. We present a detailed breakdown of test accuracy over different tasks under RewardBench with scaling in Fig. 4(b) and App.  D. As shown in Table 2, the accuracy on the Chat subset of RewardBench is low (often below random chance) with 5k examples. However, Fig. 4(a) demonstrates a significant improvement when scaling to 10k pairs, achieving very high scores (0.91) and potentially saturating this task. Compare this to Fig. 4(b) on Reasoning – a harder task – where using critiques outperforms the baseline (by more than 10 points) and where they only converge with 20k training examples. In addition, Fig. 6(b) shows that training with critiques leads to greater robustness against adversarial examples on Chat Hard, and the effect is persistent across training data sizes. Interestingly, critiques offer a more significant performance boost for tasks involving reasoning and adversarial skills, across increasing training data. We hypothesize that this is attributed to the Chain-of-Thought style feedback within critiques, which provides additional context for RMs to condition on when predicting a scalar score. We leave it as future work to explore the impact of critiques on LLM reasoning abilities when preference-tuned with such critique-enhanced RMs.
Critiques versus Knowledge Distillation
As the critiques in our experiments are generated from another model, training critiques RM can be seen as a form of knowledge distillation. To understand the benefits of critiques versus general knowledge transfer, we perform baseline studies where we directly prompt the critique generator for Likert scores (on a scale of 0 to 10), for each point-wise prompt-completion pair in the test set, and compute the proportion that is correct if the generated score is higher on the preferred completion. This setting resembles the performance of an inference-time generative RM (Li et al., 2023a) and can serve as an upper bound of the test set performance achievable from a distilled model. We call this baseline Generative Judge RM and present the results with a Command R+ critique generator (called Gen-Command R+). We further include another inference-time baseline, Gen-Command R+ (+Critiques) where we additionally prompt for a critique before outputting a score. We use the same prompt template for critiques RMs in App. A. In Table 2, we observe that while Gen-Command R+ performs strongly on Chat and comparably on Safety, the critiques RMs, across different critique generators, overwhelmingly outperform this baseline on Reasoning and Chat Hard, validating their effectiveness particularly in more challenging and adversarial prompts. For Gen-Command R+ (+Critiques), we observe that similar to training scalar RMs, conditioning on critiques during inference leads to a higher test score (with an increase of 0.3 on Reasoning) than No-Critiques, agreeing with the observations in Wang et al. (2023a). As the benefit of critiques exists in both training and inference, using critiques is not merely a form of knowledge transfer but introduces an additional Chain-of-Thought that enhances the model’s decision-making process.